# Representational Alignment Across Model Layers and Brain Regions with Hierarchical Optimal Transport

### v1

@@ Abstract
Standard representational similarity methods align each layer of a network to its best match in another independently, producing asymmetric results, lacking a global alignment score, and struggling with networks of different depths. These limitations arise from ignoring global activation structure and restricting mappings to rigid one-to-one layer correspondences. We propose Hierarchical Optimal Transport (HOT), a unified framework that jointly infers soft, globally consistent layer-to-layer couplings and neuron-level transport plans. HOT allows source neurons to distribute mass across multiple target layers while minimizing total transport cost under marginal constraints. This yields both a single alignment score for the entire network comparison and a soft transport plan that naturally handles depth mismatches through mass distribution. We evaluate HOT on vision models, large language models, and human visual cortex recordings. Across all domains, HOT matches or surpasses standard pairwise matching in alignment quality. Moreover, it reveals smooth, fine-grained hierarchical correspondences: early layers map to early layers, deeper layers maintain relative positions, and depth mismatches are resolved by distributing representations across multiple layers. These structured patterns emerge naturally from global optimization without being imposed, yet are absent in greedy layer-wise methods. HOT thus enables richer, more interpretable comparisons between representations, particularly when networks differ in architecture or depth.
@@ 1 Introduction
We propose Hierarchical Optimal Transport (HOT), a framework for globally consistent representational alignment. Optimal transport–based methods, such as Soft-Matching distance (Khosla and Williams, 2024; Khosla et al., 2024), have recently emerged as powerful metrics for comparing neural representations. Unlike metrics such as RSA, CKA, or linear predictivity—which are rotation-invariant and thus unable to capture similarities in neuron-level tuning—OT-based methods are rotation-sensitive. They explicitly match neurons based on their tuning profiles and, by relaxing hard permutation constraints into fractional couplings, can also handle layers of unequal size. This enables richer, more flexible neuron-level alignments than either rotation-invariant similarity metrics or strict permutation-based approaches. However, Soft Matching also remains limited to pairwise layer comparisons like other methods and does not capture global structure across networks. HOT fills this gap by operating hierarchically: it simultaneously infers soft neuron-to-neuron couplings within layers and a soft, globally consistent layer-to-layer coupling across the hierarchy. Rather than forcing each source layer to match exactly one target layer, HOT allows source layers to distribute their representational “mass” across multiple target layers while minimizing the total transport cost under marginal constraints; that is, each source layer must distribute exactly 100 of its mass across target layers (no information is lost), and the total mass each target layer receives from all source layers must sum to a balanced allocation (no target is over- or under-utilized). These conservation laws ensure a balanced alignment where every layer contributes meaningfully to the global correspondence, preventing any layer from being arbitrarily overweighted or ignored in the global matching. The result is a single network-level alignment score and a soft transport plan that naturally handles depth mismatches.
@@
We evaluate HOT on three diverse domains: comparisons between foundation models in vision (Vision Transformers like DINOv2 and ViT-MAE), large language models of varying scales (LLaMA, Qwen), and fMRI recordings from human visual cortex across different participants. Our key contributions are:
@@
A principled global alignment framework (theoretical): HOT jointly optimizes all layer correspondences to produce symmetric, globally consistent assignments with a single alignment score. In contrast, greedy pairwise approaches are asymmetric, can overweight certain layers while completely ignoring others, and and may spuriously treat wide layers as similar to every layer they are compared against, since their high dimensionality allows them to fit or partially overlap with many different representational subspaces, obscuring more meaningful correspondences.
@@
Natural handling of depth mismatches (theoretical): By allowing soft, many-to-many mappings between layers, HOT aligns networks of different depths without forcing inappropriate one-to-one correspondences.
@@
Rotation-invariant extension (theoretical): We propose an extension of HOT that incorporates additional orthogonal transformations, making the framework rotation-invariant. This ensures that correspondences can be recovered even when shared representational features are embedded in rotated subspaces, and yields consistently high-quality alignments.
@@
Improved alignment scores (empirical): Across domains (vision models, large language models, and brain data), HOT matches or surpasses standard pairwise methods, yielding higher alignment.
@@
Emergent hierarchical structure (empirical): Without imposing ordering constraints, HOT recovers known hierarchical organization in visual cortex data across subjects and reveals clean layer-to-layer correspondences in model–model comparisons where early layers map to early layers and deeper layers preserve their relative ordering. By contrast, greedy pairwise methods fail to reveal hierarchical structure, often leaving many layers unmatched while a single or few layers dominate the mappings.
@@
Featural distribution across depth (empirical): HOT reveals how deeper networks spread computations across multiple layers that shallower networks compress into fewer stages. Concretely, a single layer in a shallower network often distributes its mass across several neighboring layers in a deeper network, an effect that greedy pairwise methods completely miss.
@@ 2.1 Problem Setup and Existing Approaches
In what follows, we introduce a hierarchical optimal transport framework that overcomes both issues by simultaneously inferring a soft layer-to-layer coupling and per-neuron transport plans, yielding a globally balanced alignment across networks of arbitrary depths (Figure 1).
2.2 Hierarchical Optimal Transport
Our framework operates at two hierarchical levels: an inner level that aligns neurons within each layer pair, and an outer level that determines how layers should be coupled globally.
@@ 2.2.1 Inner Level: Neuron-to-Neuron Transport
For each pair of layers , we compute a cost matrix
where is the correlation distance and denotes Pearson correlation. We then solve an optimal transport problem to find a coupling that minimizes the transport cost between neurons:
where is the transportation polytope:
This constraint ensures that each source neuron distributes exactly of its mass (summing rows), and each target neuron receives exactly of the total mass (summing columns), with all entries non-negative. The coupling specifies how neurons in layer can be combined to reconstruct neurons in layer .
This inner-level optimization is precisely the Soft Matching distance (Khosla and Williams, 2024) applied between layers and : it is rotation-sensitive yet permutation-invariant, and directly measures single-neuron tuning alignment. When the layers have equal size (), the optimal solution lies at a vertex of the transportation polytope; which is a permutation matrix by the Birkhoff–von Neumann theorem (von Neumann, 1953; De Loera and Kim, 2013). However, when layer sizes differ, soft matching assigns each neuron in one layer to weighted combinations of neurons in the other via soft-assignments.
@@ 2.2.3 Reconstruction and Evaluation
The global HOT score is the average across all layers:
@@ 2.3 Rotation-Invariant Extension
Most existing metrics of representational similarity, such as RSA, CKA, and Procrustes distance, are designed to be rotation-invariant. The rationale is that when comparing population codes, we often care less about the tuning of individual neurons and more about the geometry of the representational space: distances, angles, and relative positions between stimulus responses. Two networks can encode essentially the same geometry while using different coordinate bases (for example, rotated versions of one another). Requiring strict unit-to-unit correspondence in such cases would artificially inflate dissimilarity, even though the representational geometry and information content are preserved. A rotation-invariant extension of HOT (HOT + R henceforth) therefore provides a way to capture this geometric equivalence while still enforcing a globally consistent layer- and neuron-level coupling. We introduce rotation matrices for each layer pair and minimize the reconstruction error:
@@ 2.4 Baseline Comparisons
We compare HOT against several baselines to isolate the contribution of each component:
@@
HOT Top-1 Layer Transport Plan (Single-Best OT):
@@ 3 Results
We evaluate representational similarity under the HOT metric across four distinct alignment setups: (i) large language models of different families and scales, (ii) fMRI responses from the visual cortex of four human subjects, (iii) pretrained transformer-based vision models spanning different families and scales, and (iv) cross-domain comparisons between human visual cortex and vision transformers (Appendix Section D). Across all settings, HOT matches or surpasses baseline reconstruction (prediction) scores. Importantly, the transport plans inferred by HOT naturally reveal systematic layer-wise correspondences across models and cortex, despite not being explicitly optimized for such structure; specifically, earlier layers or regions tend to align with earlier counterparts, while deeper layers or regions map to progressively higher levels; patterns that greedy pairwise methods fail to capture.
@@ 3.1 Representational Similarity Between Large Language Models
Experimental Setup. We extract layer-wise representations by averaging token activations across 2,552 prompts from the STSB dataset (May, 2021; Enevoldsen et al., 2025; Muennighoff et al., 2022). For each model, this yields a sequence of representation matrices and , corresponding to successive layers. We evaluate models of varying sizes from the LLaMA-3.2 (Grattafiori et al., 2024) and Qwen-2.5 (Qwen et al., 2025) families. Representations are compared using HOT and baseline metrics, and the resulting transport plans are analyzed. To quantify alignment quality, we reconstruct representations on a held-out validation split (20) using the learned transport maps and report the correlation with ground-truth activations, as detailed in the Methods Section2.2.3.
| Model 1 | Model 2 | HOT Metric | Random (Perm-P) | Single-Best OT | Pairwise Best OT |
@@
Results. HOT consistently achieves higher reconstruction accuracy than baseline methods. As shown in Table 1, reconstruction scores on the held-out validation set are significantly higher under HOT, indicating that the alignment plans it learns are more robust and generalizable than those obtained from pairwise baselines. This improvement stems from HOT’s (i) enforcement of global consistency across all layers and (ii) ability to distribute representational mass across multiple target layers, allowing it to recover alignments that remain hidden to pairwise methods when features from a single layer are distributed across several layers in another model. Beyond quantitative gains, HOT uncovers clear hierarchical correspondences between models. As illustrated in Figures 2 and A.1, the transport plans produced by HOT exhibit strong diagonal structure: early layers in one model align with early layers in the other, while deeper layers align with deeper layers. Such structured correspondences are absent in pairwise OT, which often yields noisy mappings. Furthermore, when comparing shallower with deeper models, HOT reveals that single layers in the shallower model distribute their mass across multiple consecutive layers in the deeper model. This soft many-to-many mapping reflects how additional depth refines and spreads computations, suggesting that representational stages compressed into one layer in a shallower model are decomposed across multiple processing steps in a deeper model—an organizational principle that greedy baselines fail to uncover.
@@ 3.2 Representational Similarity Between Visual Cortex across Subjects
We target visual cortex regions (V1–V4), treating each region as a “layer” and individual voxels as “neurons”. The visual cortex provides a strong testbed for hierarchical alignment because it is one of the best-characterized cortical systems: early areas (V1, V2) are known to encode low-level features such as orientation and contrast, while higher areas (V3, V4) encode progressively more complex shapes and object features (Hubel and Wiesel, 1968; Pasupathy and Connor, 2001; Desimone and Schein, 1987; Desimone et al., 1984). This hierarchical progression is well-established across individuals, so correspondences between homologous regions are strongly expected. HOT is applied to align responses across subjects by generating transport maps between regions. To evaluate alignment quality, we reconstruct held-out responses (20 validation split) and report correlations with ground-truth activity. We repeat this evaluation across 5 random train-validation splits to ensure robustness.
| Model 1 | Model 2 | HOT Metric | Random (Perm-P) | Single-Best OT | Pairwise Best OT |
| Subject A | Subject B | 0.308 | |||
| Subject A | Subject C | 0.232 | |||
| Subject A | Subject D | 0.253 | |||
| Subject B | Subject C | 0.237 | 0.237 | 0.237 | |
| Subject C | Subject D | 0.239 | |||
| Subject B | Subject D | 0.254 |
Results. As shown in Tables 2 and B.1, HOT achieves reconstruction scores comparable to pairwise OT, with only a marginal decrease in correlation. More importantly, Figures 3 and B.1 show that the transport maps inferred by HOT recover the expected cross-subject correspondences: cortical regions in one subject consistently align to the same regions in another. In contrast, pairwise OT does not produce such structured mappings in any subject pair. This indicates that HOT’s global alignment scheme captures region-to-region correspondences that pairwise methods miss. To rule out the possibility that this effect stems from the OT optimization framework itself rather than from the layer-wise setup, we additionally perform pairwise linear predictivity optimization. As shown in Figure B.2, the best-matching layers identified by linear predictivity also do not align corresponding regions, further confirming that the hierarchical alignment mechanism in HOT is critical for recovering robust and generalizable correspondences. Overall, these results demonstrate that HOT yields transport plans that are both interpretable and biologically meaningful, while maintaining reconstruction performance on par with baseline methods.
@@ 3.3 Representation Similarity between Vision Models
Experimental Setup. We extract layer-wise representations from Vision Transformers by averaging patch activations for each input image. We use 20,000 images randomly sampled from the ImageNet validation set (Deng et al., 2009; Russakovsky et al., 2015), sampled to ensure a uniform coverage across all classes in Imagenet. For each model, this yields a sequence of representation matrices that serve as inputs to the HOT framework. We evaluate two families of pretrained vision transformers—DINOv2 and ViT-MAE across multiple model scales. Alignment quality is quantified by reconstructing representations on a held-out validation split (20) using the learned transport plans, with reconstruction–ground truth correlation as the metric.
Prior work has shown that the residual stream in Transformers lacks privileged axes and is invariant up to rotations of its basis (Khosla et al., 2024). Since HOT, like other OT-based methods, is rotation-sensitive, we additionally evaluate a rotation-augmented variant (HOT+R) and its baselines (see Methods Section 2.3). We restrict this analysis to Vision Transformers, where rotational invariances are especially relevant and where the computational cost of HOT+R remains tractable. For large language models and fMRI data, the added optimization over rotations was computationally prohibitive, so we report only the rotation-sensitive results in those domains. In HOT+R, the learned rotation matrices are incorporated into both transport optimization and evaluation, ensuring that geometric equivalences induced by rotations are properly captured.
| Model 1 | Model 2 | HOT Metric | Pairwise Best OT | HOT + R | Pairwise Best + R |
@@
Results. Table 3 shows that, on its own, vanilla HOT does not consistently outperform pairwise OT in reconstruction accuracy. Consistent with this, the transport plans inferred by HOT (Figure 4 and Figures C.1 - C.10) only partially reveal layer-wise correspondences: in some cases, clear diagonal structure emerges, but in others the mappings are noisier. By contrast, the rotation-augmented variant (HOT+R) yields substantially higher reconstruction scores surpassing both vanilla HOT and pairwise vanilla and rotational baselines (Tables 3, C.1 and C.2). Importantly, the transport plans produced by HOT+R consistently exhibit strong hierarchical correspondences, recovering clean layer-to-layer alignment even in settings where vanilla HOT fails to do so. These findings indicate that incorporating rotation into the OT framework not only improves quantitative alignment quality but also produces more generalizable and interpretable mappings, particularly in domains like Vision Transformers where representations are known to be rotation-invariant.

### v2

# Representational Alignment Across Model Layers and Brain Regions with Multi-Level Optimal Transport

@@ Abstract
Standard representational similarity methods align each layer of a network to its best match in another independently, producing asymmetric results, lacking a global alignment score, and struggling with networks of different depths. These limitations arise from ignoring global activation structure and restricting mappings to rigid one-to-one layer correspondences. We propose Multi-Level Optimal Transport (MOT), a unified framework that jointly infers soft, globally consistent layer-to-layer couplings and neuron-level transport plans. MOT allows source neurons to distribute mass across multiple target layers while minimizing total transport cost under marginal constraints. This yields both a single alignment score for the entire network comparison and a soft transport plan that naturally handles depth mismatches through mass distribution. We evaluate MOT on vision models, large language models, and human visual cortex recordings. Across all domains, MOT matches or surpasses standard pairwise matching in alignment quality. Moreover, it reveals smooth, fine-grained hierarchical correspondences: early layers map to early layers, deeper layers maintain relative positions, and depth mismatches are resolved by distributing representations across multiple layers. These structured patterns emerge naturally from global optimization without being imposed, yet are absent in greedy layer-wise methods. MOT thus enables richer, more interpretable comparisons between representations, particularly when networks differ in architecture or depth. We further extend our method to a three-level MOT framework, providing a proof-of-concept alignment of two networks across their training trajectories and demonstrating that MOT uncovers checkpoint-wise correspondences missed by greedy layer-wise matching.
@@ 1 Introduction
We propose Multi-Level Optimal Transport (MOT), a framework for globally consistent representational alignment. Optimal transport–based methods, such as Soft-Matching distance (Khosla and Williams, 2024; Khosla et al., 2024), have recently emerged as powerful metrics for comparing neural representations. Unlike metrics such as RSA, CKA, or linear predictivity—which are rotation-invariant and thus unable to capture similarities in neuron-level tuning—OT-based methods are rotation-sensitive. They explicitly match neurons based on their tuning profiles and, by relaxing hard permutation constraints into fractional couplings, can also handle layers of unequal size. This enables richer, more flexible neuron-level alignments than either rotation-invariant similarity metrics or strict permutation-based approaches. However, Soft Matching also remains limited to pairwise layer comparisons like other methods and does not capture global structure across networks. MOT fills this gap by operating across multiple levels: it simultaneously infers soft neuron-to-neuron couplings within layers and a soft, globally consistent layer-to-layer coupling across the two levels. Rather than forcing each source layer to match exactly one target layer, MOT allows source layers to distribute their representational “mass” across multiple target layers while minimizing the total transport cost under marginal constraints; that is, each source layer must distribute exactly 100 of its mass across target layers (no information is lost), and the total mass each target layer receives from all source layers must sum to a balanced allocation (no target is over- or under-utilized). These conservation laws ensure a balanced alignment where every layer contributes meaningfully to the global correspondence, preventing any layer from being arbitrarily overweighted or ignored in the global matching. The result is a single network-level alignment score and a soft transport plan that naturally handles depth mismatches.
@@
We evaluate MOT on three diverse domains: comparisons between foundation models in vision (Vision Transformers like DINOv2 and ViT-MAE), large language models of varying scales (LLaMA, Qwen), and fMRI recordings from human visual cortex across different participants. Our key contributions are:
@@
A principled global alignment framework (theoretical): MOT jointly optimizes all layer correspondences to produce symmetric, globally consistent assignments with a single alignment score. In contrast, greedy pairwise approaches are asymmetric, can overweight certain layers while completely ignoring others, and may spuriously treat wide layers as similar to every layer they are compared against, since their high dimensionality allows them to fit or partially overlap with many different representational subspaces, obscuring more meaningful correspondences.
@@
Natural handling of depth mismatches (theoretical): By allowing soft, many-to-many mappings between layers, MOT aligns networks of different depths without forcing inappropriate one-to-one correspondences.
@@
Rotation-invariant extension (theoretical): We propose an extension of MOT that incorporates additional orthogonal transformations, making the framework rotation-invariant. This ensures that correspondences can be recovered even when shared representational features are embedded in rotated subspaces, and yields consistently high-quality alignments.
@@
Improved alignment scores (empirical): Across domains (vision models, large language models, and brain data), MOT matches or surpasses standard pairwise methods, yielding higher alignment.
@@
Emergent hierarchical structure (empirical): Without imposing ordering constraints, MOT recovers known hierarchical organization in visual cortex data across subjects and reveals clean layer-to-layer correspondences in model–model comparisons where early layers map to early layers and deeper layers preserve their relative ordering. By contrast, greedy pairwise methods fail to reveal hierarchical structure, often leaving many layers unmatched while a single or few layers dominate the mappings.
@@
Featural distribution across depth (empirical): MOT reveals how deeper networks spread computations across multiple layers that shallower networks compress into fewer stages. Concretely, a single layer in a shallower network often distributes its mass across several neighboring layers in a deeper network, an effect that greedy pairwise methods completely miss.
@@ 2.1 Problem Setup and Existing Approaches
2.2 Multi-Level Optimal Transport
Our framework operates at two levels: an inner level that aligns neurons within each layer pair, and an outer level that determines how layers should be coupled globally.
@@ 2.2.1 Inner Level: Neuron-to-Neuron Transport
For each candidate layer pair , MOT computes a neuron-level alignment cost by applying the soft-matching OT (Appendix A for more details) formulation to the tuning functions in and . We first construct a pairwise dissimilarity matrix
where is the chosen tuning-based dissimilarity (here, correlation distance). The inner OT objective then computes
yielding a soft coupling that specifies how strongly each neuron in layer corresponds to neurons in layer .
Because lies in the transportation polytope, its row and column marginals ensure that each source neuron distributes its representational mass uniformly and each target neuron receives its appropriate share. When the two layers have the same width (), the optimal solution reduces to a permutation (by the Birkhoff–von Neumann theorem (von Neumann, 1953; De Loera and Kim, 2013)); when they differ, soft assignments naturally emerge, allowing neurons in one layer to map to mixtures of neurons in the other. The resulting inner OT cost serves as the layer-to-layer dissimilarity used by the outer level of MOT.
@@ 2.2.3 Reconstruction and Evaluation
The global MOT score is the average across all layers:
@@ 2.3 Rotation-Invariant Extension
Most existing metrics of representational similarity, such as RSA, CKA, and Procrustes distance, are designed to be rotation-invariant. The rationale is that when comparing population codes, we often care less about the tuning of individual neurons and more about the geometry of the representational space: distances, angles, and relative positions between stimulus responses. Two networks can encode essentially the same geometry while using different coordinate bases (for example, rotated versions of one another). Requiring strict unit-to-unit correspondence in such cases would artificially inflate dissimilarity, even though the representational geometry and information content are preserved. A rotation-invariant extension of MOT (MOT + R henceforth) therefore provides a way to capture this geometric equivalence while still enforcing a globally consistent layer- and neuron-level coupling. We introduce rotation matrices for each layer pair and minimize the reconstruction error:
@@ 2.4 Baseline Comparisons
We compare MOT against several baselines to isolate the contribution of each component:
@@
MOT Top-1 Layer Transport Plan (Single-Best OT):
@@ 3 Results
We evaluate representational similarity under the MOT metric across four distinct alignment setups: (i) large language models of different families and scales, (ii) fMRI responses from the visual cortex of four human subjects, (iii) pretrained transformer-based vision models spanning different families and scales, (iv) cross-domain comparisons between human visual cortex and vision transformers (Appendix Section E). Across all settings, MOT matches or surpasses baseline reconstruction (prediction) scores. Importantly, the transport plans inferred by MOT naturally reveal systematic layer-wise correspondences across models and cortex, despite not being explicitly optimized for such structure; specifically, earlier layers or regions tend to align with earlier counterparts, while deeper layers or regions map to progressively higher levels; patterns that greedy pairwise methods fail to capture. Furthermore, we demonstrate a proof-of-concept extension of MOT to a Three-level MOT framework by aligning two networks across their training trajectories.
@@ 3.1 Representational Similarity Between Large Language Models
Experimental Setup. We extract layer-wise representations by averaging token activations across 2,552 prompts from the STSB dataset (May, 2021; Enevoldsen et al., 2025; Muennighoff et al., 2022). For each model, this yields a sequence of representation matrices and , corresponding to successive layers. We evaluate models of varying sizes from the LLaMA-3.2 (Grattafiori et al., 2024) and Qwen-2.5 (Qwen et al., 2025) families. Representations are compared using MOT and baseline metrics, and the resulting transport plans are analyzed. To quantify alignment quality, we reconstruct representations on a held-out validation split (20) using the learned transport maps and report the correlation with ground-truth activations, as detailed in the Methods Section2.2.3.
| Model 1 | Model 2 | MOT Metric | Random (Perm-P) | Single-Best OT | Pairwise Best OT |
@@
Results. MOT consistently achieves higher reconstruction accuracy than baseline methods. As shown in Table 1, reconstruction scores on the held-out validation set are significantly higher under MOT, indicating that the alignment plans it learns are more robust and generalizable than those obtained from pairwise baselines. This improvement stems from MOT’s (i) enforcement of global consistency across all layers and (ii) ability to distribute representational mass across multiple target layers, allowing it to recover alignments that remain hidden to pairwise methods when features from a single layer are distributed across several layers in another model. Beyond quantitative gains, MOT uncovers clear hierarchical correspondences between models. As illustrated in Figures 2 and B.1, the transport plans produced by MOT exhibit strong diagonal structure: early layers in one model align with early layers in the other, while deeper layers align with deeper layers. Such structured correspondences are absent in pairwise OT, which often yields noisy mappings. Furthermore, when comparing shallower with deeper models, MOT reveals that single layers in the shallower model distribute their mass across multiple consecutive layers in the deeper model. This soft many-to-many mapping reflects how additional depth refines and spreads computations, suggesting that representational stages compressed into one layer in a shallower model are decomposed across multiple processing steps in a deeper model—an organizational principle that greedy baselines fail to uncover.
To rule out the possibility that this effect stems from the multi-level formulation and not from OT, we also perform pairwise linear predictivity and RSA analyses. As shown in Figures H.1 and I.1, these methods also fail to reveal hierarchial correspondence.
@@ 3.2 Representational Similarity Between Visual Cortex across Subjects
We target visual cortex regions (V1–V4, Lateral, Dorsal, Ventral), treating each region as a “layer” and individual voxels as “neurons”. The visual cortex provides a strong testbed for multi-level alignment because it is one of the best-characterized cortical systems: early areas (V1, V2) are known to encode low-level features such as orientation and contrast, while higher areas (V3, V4) encode progressively more complex shapes and object features (Hubel and Wiesel, 1968; Pasupathy and Connor, 2001; Desimone and Schein, 1987; Desimone et al., 1984). This hierarchical progression is well-established across individuals, so correspondences between homologous regions are strongly expected. MOT is applied to align responses across subjects by generating transport maps between regions. To evaluate alignment quality, we reconstruct held-out responses (20 validation split) and report correlations with ground-truth activity. We repeat this evaluation across 5 random train-validation splits to ensure robustness.
| Model 1 | Model 2 | MOT Metric | Random (Perm-P) | Single-Best OT | Pairwise Best OT |
| Subject A | Subject B | 0.245 | |||
| Subject A | Subject C | 0.202 | |||
| Subject A | Subject D | 0.201 | |||
| Subject B | Subject C | 0.212 | 0.212 | 0.212 | |
| Subject C | Subject D | 0.199 | |||
| Subject B | Subject D | 0.204 |
Results. As shown in Tables 2 and C.1, MOT achieves reconstruction scores comparable to pairwise OT, with only a marginal decrease in correlation. More importantly, Figures 3 and C.1 show that the transport maps inferred by MOT recover the expected cross-subject correspondences: cortical regions in one subject consistently align to the same regions in another. In contrast, pairwise OT does not produce such structured mappings in any subject pair. This indicates that MOT’s global alignment scheme captures region-to-region correspondences that pairwise methods miss. We further assess pairwise linear predictivity and RSA, with the results as shown in Figures H.1 and I.1. The layers selected as best-matching by these approaches fail to align corresponding regions as well, reinforcing that the multi-level alignment mechanism in MOT is crucial for obtaining robust and generalizable correspondences. Overall, these results demonstrate that MOT yields transport plans that are both interpretable and biologically meaningful, while maintaining reconstruction performance on par with baseline methods.
@@ 3.3 Representation Similarity between Vision Models
Experimental Setup. We extract layer-wise representations from Vision Transformers by averaging patch activations for each input image. We use 20,000 images randomly sampled from the ImageNet validation set (Deng et al., 2009; Russakovsky et al., 2015), sampled to ensure a uniform coverage across all classes in Imagenet. For each model, this yields a sequence of representation matrices that serve as inputs to the MOT framework. We evaluate two families of pretrained vision transformers—DINOv2 and ViT-MAE across multiple model scales. Alignment quality is quantified by reconstructing representations on a held-out validation split (20) using the learned transport plans, with reconstruction–ground truth correlation as the metric.
Prior work has shown that the residual stream in Transformers lacks privileged axes and is invariant up to rotations of its basis (Khosla et al., 2024). Since MOT, like other OT-based methods, is rotation-sensitive, we additionally evaluate a rotation-augmented variant (MOT+R) and its baselines (see Methods Section 2.3). We restrict this analysis to Vision Transformers, where rotational invariances are especially relevant and where the computational cost of MOT+R remains tractable. For large language models and fMRI data, the added optimization over rotations was computationally prohibitive, so we report only the rotation-sensitive results in those domains. In MOT+R, the learned rotation matrices are incorporated into both transport optimization and evaluation, ensuring that geometric equivalences induced by rotations are properly captured.
| Model 1 | Model 2 | MOT Metric | Pairwise Best OT | MOT + R | Pairwise Best + R |
@@
Results. Table 3 shows that, on its own, vanilla MOT does not consistently outperform pairwise OT in reconstruction accuracy. The transport plans inferred by MOT (Figure 4 and Figures D.1 - D.10) only partially reveal layer-wise correspondences: in some cases, clear diagonal structure emerges, but in others the mappings are noisier. By contrast, the rotation-augmented variant (MOT+R) yields substantially higher reconstruction scores surpassing both vanilla MOT and pairwise vanilla and rotational baselines (Tables 3, D.1 and D.2). Importantly, the transport plans produced by MOT+R consistently exhibit strong hierarchical correspondences, recovering clean layer-to-layer alignment even in settings where vanilla MOT fails to do so. These findings indicate that incorporating rotation into the OT framework not only improves quantitative alignment quality but also produces more generalizable and interpretable mappings, particularly in domains like Vision Transformers where representations are known to be rotation-invariant.
3.4 Three-Level MOT for Aligning Training Trajectories
Formulation of Three-Level MOT. We extend the two-level MOT formulation described above to an additional level over checkpoints. Let the first model have checkpoints and the second , with checkpoint- activations and checkpoint- activations .
For each checkpoint pair , we treat and as two networks and run the same two-level MOT procedure as above, yielding neuron- and layer-level couplings together with a single scalar cost
where denotes the two-level objective (inner neuron OT followed by outer layer OT). This defines a checkpoint-level cost matrix . To obtain a globally consistent alignment between the two training trajectories, we then solve a third OT problem over checkpoints, reusing the same uniform-marginal formulation as before:
and can be interpreted as a soft assignment between checkpoint in the first model and checkpoint in the second. For comparison, we also construct a simple pairwise-best checkpoint map that ignores the global OT constraint and matches each checkpoint independently to its lowest-cost partner.
Experimental Setup and Results. We perform three-level OT (and the corresponding pairwise baseline) as a proof of concept for 154 checkpoints of Pythia-14M models with seed 1 and seed 2 (Biderman et al. (2023)), where the 154 checkpoints correspond to steps 0 (initialization), 1, 2, 4, 8, 16, 32, 64, 128, 256, 512, 1000, and then every 1,000 subsequent steps. The layer-wise representations for each checkpoint are collected across 2,552 prompts from the STSB dataset (May, 2021; Enevoldsen et al., 2025; Muennighoff et al., 2022). We plot the checkpoint-level transport maps generated by three-level MOT and by the pairwise-best alignment in Figure 5. MOT reveals a coherent checkpoint-wise correspondence between the two training trajectories, whereas pairwise matching fails to do so. This demonstrates a proof-of-concept method for aligning entire training trajectories of two models. In principle, the MOT framework can be generalized to any number of levels, depending on the type of alignment setting being studied.